OpenStudy (anonymous):

given the following unbalanced equation: 8Al + 3Fe3O4 ---> 9Fe + 4Al2O3 how much Fe and Al2O3 will pe produced?

OpenStudy (anonymous):

if 25 g of Al and 700 g of Fe3O4 is reacted

OpenStudy (anonymous):

The reason I left this out is because I don't want an answer. I want to know how to do it.

mathslover (mathslover):

@vivisong10 , First of all , Welcome to OpenStudy! Thanks for asking a question here. Now, please recheck your question. The equation provided there : \(\bf{8Al + 3{Fe}_3 O_4 \rightarrow 9Fe + 4{Al}_2 O_3 }\) is actually balanced. You can check it like this : \(\begin{array}{|c|c|c|} \hline \text{} &\textbf{Left Hand Side (No. of atoms)} & \textbf{Right Hand Side (No. of atoms)} \\ \hline \mathsf{Aluminium (Al)} & \color{green}{8*1 =8} & \color{blue}{4*2 = 8} \\ \hline \mathsf{Iron (Fe)} & \color{green}{3*3 = 9} & \color{blue}{9 * 1 = 9 } \\ \hline \mathsf{Oxygen (O) } & \color{green}{3*4 = 12} & \color{blue}{4 * 3 = 12} \\ \hline \end{array} \) You can notice from the above table, that the no. of atoms of each element are same in both sides (Left hand side as well as in Right Hand Side) . Thus, the equation is balanced.

mathslover (mathslover):

Notice that 9Fe and \(4{Fe}_3 O_4\) are produced , which means 9 moles of Iron (Fe) and 4 moles of \(Fe_3O_4\) are produced in the equation. I assume, that you will have to balance this equation first and then find the above answer : \(\bf{Al + Fe_3O_4 \rightarrow Fe + Al_2O_3}\) Hint : To balance the above equation, try to equal the no. of atoms of each element on both sides. I hope, this resolves your problem. let us know if you liked it or you have any problem related to it. Come back soon on OpenStudy and answer other questions too, that will increase your knowledge also. You can ask questions too , but don't forget to give medal to those who helped you! :) Good Luck!